# Practical Aspects on Solving Differential Equations Using Deep Learning: A Primer

### v1

@@ 3.1 Feed-forward neural network
Let’s delve into the practical application of feed-forward neural networks, also known as multi-layer perceptron (MLP). These networks are not just theoretical concepts but powerful tools that map some input ${\bf x}\in\mathbb{R}^{m}$ to some output ${\bf y}\in\mathbb{R}^{p}$. In simpler terms, a neural network defines a mapping ${\bf y}=\mathcal{F}(\boldsymbol{\theta};{\bf x})$ between an input and an output, making it a key component in the field of machine learning and artificial intelligence.
@@
Figure 2 A shows a typical graph of a feed-forward neural network, where we see the input (green) and output (red) layers as well as the hidden (gray boxes) layers. We can also see the connectivity matrix from the input to the first hidden layer. We omit the rest of the connections due to simplicity. The feed-forward neural network in Figure 2 A has three hidden layers, one input, and one output layer. Hence, we can write the operations it performs as ${\bf y}=f({\bf x})=f^{(4)}\circ f^{(3)}\circ f^{(2)}\circ f^{(1)}({\bf x})=f^{(4)}(f^{(3)}(f^{(2)}(f^{(1)}({\bf x}))))$, where $f_{1}$ is the mapping from the input to the first hidden layer, $f_{2}$ is the mapping from the first hidden to the second and so on so forth. If we are to expand the composition by using affine transformations, we have ${\bf y}=f({\bf x})={\bf W}_{4}\cdot(g_{3}({\bf W}_{3}\cdot g_{2}({\bf W}_{2}\cdot g_{1}({\bf W}_{1}\dot{\bf x}+{\bf b}_{1})+{\bf b}_{2})+{\bf b}_{3})+{\bf b}_{4}$, where $g_{l}$ represents the activation functions in each layer $l$. Notice that the output layer $L$ has no activation function.
@@ 3.5 Backpropagation, Optimizers and Automatic Differentiation
where $\eta$ is the learning rate. When we use mini-batches, we use a modified version of the GD called Stochastic Gradient Descend (SGD), which looks like $\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};{\bf x},{\bf y})=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\boldsymbol{\theta}}L_{i}(\boldsymbol{\theta};{\bf x}_{i},{\bf y}_{i})$, where $n$ is the size of the batch and $L_{i}$ is the loss per training example [55, 25].
The SGD or any other optimizer we might use requires an extensive computation of partial derivatives (gradients) with respect to the loss function. All modern machine learning and deep learning frameworks implement the backprop algorithm using Automatic Differentiation (AD). AD is a collection of methods to compute and evaluate partial derivatives of functions specified by computer programs [3]. AD should not be confused with numerical or symbolic differentiation. AD builds a computational graph of all the operations within a neural network based on primitive (basic) operations and elementary functions. Applying the chain rule over and over on the computational graph can evaluate all the partial derivatives needed by the SGD. For more details on optimizers, backprop, and AD, see [25, 11].
@@
Algorithm 1 summarizes the learning steps, and code listing 4 shows the backbone of training a neural network in Pytorch. In general, we first instantiate our neural network class (line $2$ in code listing 4, and then we specify the optimizer (in this case, at line $4$, we use the Adam optimizer [40]). In this work, we define custom loss functions per equation; we will see in Section 4.1 how to implement a custom loss function. Then, for each iteration (line $7$), we sample our independent variables (e.g., time $t$, space $x$, or both) based on predefined distributions. Once we get the samples, we zero the grads (line $10$), then we approximate the solution yhat using the neural network and the independent variables. Finally, we pass the approximation to the loss function, do a backward step (lines $13$-$18$), and finally minimize it (line $21$).
@@ 3.7 Batch Normalization
In the previous section, we discussed the vanishing and exploding gradients problem that might arise when training a neural network. Generalization errors are another potential problem we face while training a neural network. The generalization error is the error a neural network commits when operating on data it has not previously trained on [25]. Typically, we want a neural network to perform well even on data that has not been seen before (provided that the data still comes from the same distribution as the training data set). When a neural network performs well even on unseen data, it generalizes well. Usually, we split the data set into three disjoint subsets: training, test, and validation data sets, and we evaluate the generalization error on the test data set. The problem arises when we end up with a simple neural network that cannot “learn” the primary trends and statistics of the data and has a significant generalization error. In this case, we say that the neural network is underfitting the data. On the other hand, we may end up with a complex neural network, or we do not have enough training data to train the network, leading to a model that “learns” the training data too well without generalizing [11]. In this case, we say that the neural network is overfitting the data.
@@ 3.8 Universal Approximation Theorem
Figure 3: Universal Approximation Theorem. A neural network with one hidden layer, three hidden units, and a $\tanh$ activation function approximates the function $f(x)=\sin(3x)$ in the interval $[-1,1]$. Fifty samples of the function $f(x)$ used to train the neural network are shown here as black dots. The orange line indicates the approximation, $\hat{f}(x)$ provided by the neural network.
@@
In simple words, a feed-forward neural network with a single hidden layer followed by a nonlinear activation function such as a sigmoid ($\sigma(x)=\frac{1}{1+\exp(-x)}$) and a linear output layer can approximate a continuous function defined on a compact subset of $\mathbb{R}^{q}$. Although the theorem states that it is possible to approximate any such function, it becomes apparent in practice that it needs to discuss the number of hidden units we will have to use to approximate a function of interest. However, deep feed-forward neural networks seem to work in practice and approximate functions defined on compact subsets of $\mathbb{R}^{q}$. Furthermore, a variation of the Universal Approximation theorem regarding the first derivatives of continuous functions has been proposed by Hornik et al. [36].
@@ 4.1 Deep Galerkin Method
$\begin{split}\mathcal{J}(\mathcal{F}(\boldsymbol{\theta};{\bf x},t))&=\underbrace{||(\partial_{t}+\mathcal{T})\mathcal{F}(\boldsymbol{\theta};{\bf x},t)||_{[0,T]\times\Omega,\nu_{1}}^{2}}_{\text{Differential Operator}}+\underbrace{||\mathcal{F}(\boldsymbol{\theta};{\bf x},t)-g({\bf x},t)||_{[0,T]\times\partial\Omega,\nu_{2}}^{2}}_{\text{Boundary Conditions}}+\underbrace{||\mathcal{F}(\boldsymbol{\theta};{\bf x},0)-y_{0}({\bf x})||_{\Omega,\nu_{3}}^{2}}_{\text{Initial Conditions}},\end{split}$ (14)
@@
$\mathcal{L}(\boldsymbol{\theta})$ $=\Big{(}\partial_{t}\mathcal{F}(\boldsymbol{\theta};{\bf x}_{\Omega},t_{\Omega})+\mathcal{T}\mathcal{F}(\boldsymbol{\theta};{\bf x},t_{\Omega})\Big{)}^{2}+\Big{(}\mathcal{F}(\boldsymbol{\theta};{\bf x}_{\partial\Omega},t_{\partial\Omega})-g({\bf x}_{\partial\Omega},t_{\partial\Omega})\Big{)}^{2}+\Big{(}\mathcal{F}(\boldsymbol{\theta};{\bf x}_{t_{0}},0)-y_{0}({\bf x}_{t_{0}})\Big{)}^{2},$ (15)
@@ 15 L_init = ((y0 - y_ic)**2)
Listing 8: Loss function for the exponential decay ODE.

### v2

@@ 3.1 Feed-forward neural network
Let’s delve into the practical application of feed-forward neural networks, also known as multi-layer perceptron (MLP). These networks are not just theoretical concepts but powerful tools that map some input ${\bf x}\in\mathbb{R}^{m}$ to some output ${\bf y}\in\mathbb{R}^{p}$. In simpler terms, a neural network defines a mapping ${\bf y}=\mathcal{F}({\bf x};\boldsymbol{\theta})$ between an input and an output, making it a key component in the field of machine learning and artificial intelligence.
@@
Figure 2 A shows a typical graph of a feed-forward neural network, where we see the input (green) and output (red) layers as well as the hidden (gray boxes) layers. We can also see the connectivity matrix from the input to the first hidden layer. We omit the rest of the connections due to simplicity. The feed-forward neural network in Figure 2 A has three hidden layers, one input, and one output layer. Hence, we can write the operations it performs as ${\bf y}=\mathcal{F}({\bf x};\boldsymbol{\theta})=f^{(4)}\circ f^{(3)}\circ f^{(2)}\circ f^{(1)}({\bf x})=f^{(4)}(f^{(3)}(f^{(2)}(f^{(1)}({\bf x}))))$, where $f_{1}$ is the mapping from the input to the first hidden layer, $f_{2}$ is the mapping from the first hidden to the second and so on so forth. If we are to expand the composition by using affine transformations, we have ${\bf y}=\mathcal{F}({\bf x};\boldsymbol{\theta})={\bf W}_{4}\cdot(g_{3}({\bf W}_{3}\cdot g_{2}({\bf W}_{2}\cdot g_{1}({\bf W}_{1}\dot{\bf x}+{\bf b}_{1})+{\bf b}_{2})+{\bf b}_{3})+{\bf b}_{4}$, where $g_{l}$ represents the activation functions in each layer $l$. Notice that the output layer $L$ has no activation function.
@@ 3.5 Backpropagation, Optimizers and Automatic Differentiation
where $\eta$ is the learning rate. When the loss (error) function is defined with respect to an entire training data set, we must use all the available training data at each iteration of the GD. Due to their high computational cost, these so-called batch methods are not very useful when training deep neural networks. To mitigate that problem, we assume that the gradient is an expectation, and thus, by sampling our training dataset using small subsets of data, we can reasonably approximate the gradient. Hence, we use a modified version of the GD called Stochastic Gradient Descend (SGD), which looks like $\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};{\bf x},{\bf y})=\frac{1}{n}\nabla_{\boldsymbol{\theta}}\sum_{i=1}^{n}L_{i}(\boldsymbol{\theta};{\bf x}_{i},{\bf y}_{i})$, where $n$ is the size of a mini-batch (small sample uniformly drawn from the training data set) and $L_{i}$ is the loss per training example [56, 25, 7].
The SGD or any other optimizer we might use requires an extensive computation of partial derivatives (gradients) with respect to the loss function. All modern machine learning and deep learning frameworks implement the backprop algorithm using Automatic Differentiation (AD). AD is a collection of methods to compute and evaluate partial derivatives of functions specified by computer programs [3]. AD should not be confused with numerical or symbolic differentiation. AD builds a computational graph of all the operations within a neural network based on primitive (elementary) arithmetic operations and elementary functions (e.g., $\exp$, $\log$, $\cos$). Applying the chain rule over and over on the computational graph can evaluate all the partial derivatives needed by the SGD. For more details on optimizers, backprop, and AD, see [25, 11, 7].
@@
Algorithm 1 summarizes the learning steps, and code listing 4 shows the backbone of training a neural network in Pytorch. In general, we first instantiate our neural network class (line $2$ in code listing 4), then we specify the optimizer (in this case, at line $4$ we use the Adam optimizer [40]), and the criterion, which is the loss function we are interested in minimizing. In this work, we define custom loss functions per equation; we will see in Section 4.1 how to implement a custom loss function. Moreover, we might want to initialize a learning rate scheduler, which will change (usually decrease) the learning rate based on some linear, exponential, sinusoidal, or other relation [^6] The reader can find a list of learning rate schedulers already implemented in PyTorch here: https://pytorch.org/docs/stable/optim.html#how-to-adjust-learning-rate. Then, for each iteration (line $7$), we sample our independent variables (e.g., time $t$, space $x$, or both) based on predefined distributions. Once we get the samples, we zero the grads (line $10$), then we approximate the solution yhat using the neural network and the independent variables. Finally, we pass the approximation to the loss function, do a backward step (lines $13$-$18$), and finally minimize it (line $21$).
@@ 3.7 Batch Normalization
In the previous section, we discussed the vanishing and exploding gradients problem that might arise when training a neural network. Generalization errors are another potential problem we face while training a neural network. The generalization error is the error a neural network commits when operating on data it has not previously trained on [25]. Typically, we want a neural network to perform well even on data it has not seen before (provided that the data still comes from the same distribution as the training data set). When a neural network performs well even on unseen data, it generalizes well. Usually, we split the data set into three disjoint subsets: training, test, and validation data sets, and we evaluate the generalization error on the test data set. The problem arises when we end up with a simple neural network that cannot “learn” the primary trends and statistics of the data and has a significant generalization error. In this case, we say that the neural network is underfitting the data. On the other hand, we may end up with a complex neural network, or we do not have enough training data to train the network, leading to a model that “learns” the training data too well without generalizing [11]. In this case, we say that the neural network is overfitting the data.
@@ 3.8 Universal Approximation Theorem
Figure 3: Universal Approximation Theorem. A neural network with one hidden layer, three hidden units, and a $\tanh$ activation function approximates the function $f(x)=\sin(3x)$ in the interval $[-1,1]$. Fifty samples of the function $f(x)$ used to train the neural network are shown here as black dots. The orange line indicates the approximation, $\hat{f}(x)$, provided by the neural network.
@@
In simple words, a feed-forward neural network with a single hidden layer followed by a nonlinear activation function such as a sigmoid ($\sigma(x)=\frac{1}{1+\exp(-x)}$) and a linear output layer can approximate a continuous function defined on a compact subset of $\mathbb{R}^{q}$. Although the theorem states that it is possible to approximate any such function, it becomes apparent in practice that it needs to discuss the number of hidden units we will have to use to approximate a function of interest. However, deep feed-forward neural networks seem to work in practice and approximate functions defined on compact subsets of $\mathbb{R}^{q}$. Furthermore, a variation of the Universal Approximation theorem (UAT) regarding non-polynomial functions, including the ReLU functions, has been proposed by Leshno et al. [46]. In addition, a variation to the UAT involving the first derivatives of continuous functions has been suggested by Hornik et al. [36].
@@ 4.1 Deep Galerkin Method
$\begin{split}\mathcal{J}(\mathcal{F}({\bf x},t;\boldsymbol{\theta}))&=\underbrace{||(\partial_{t}+\mathcal{T})\mathcal{F}({\bf x},t;\boldsymbol{\theta})||_{[0,T]\times\Omega,\nu_{1}}^{2}}_{\text{Differential Operator}}+\underbrace{||\mathcal{F}({\bf x},t;\boldsymbol{\theta})-g({\bf x},t)||_{[0,T]\times\partial\Omega,\nu_{2}}^{2}}_{\text{Boundary Conditions}}+\underbrace{||\mathcal{F}({\bf x},0;\boldsymbol{\theta})-y_{0}({\bf x})||_{\Omega,\nu_{3}}^{2}}_{\text{Initial Conditions}},\end{split}$ (14)
@@
$\mathcal{L}(\boldsymbol{\theta})$ $=\Big{(}\partial_{t}\mathcal{F}({\bf x}_{\Omega},t_{\Omega};\boldsymbol{\theta})+\mathcal{T}\mathcal{F}({\bf x},t_{\Omega};\boldsymbol{\theta})\Big{)}^{2}+\Big{(}\mathcal{F}({\bf x}_{\partial\Omega},t_{\partial\Omega};\boldsymbol{\theta})-g({\bf x}_{\partial\Omega},t_{\partial\Omega})\Big{)}^{2}+\Big{(}\mathcal{F}({\bf x}_{t_{0}},0;\boldsymbol{\theta})-y_{0}({\bf x}_{t_{0}})\Big{)}^{2},$ (15)
@@ 15 L_init = ((y0 - y_ic)**2)
Listing 8: Loss function for the FitzHugh-Nagumo system.